AN EXPERIMENTAL DETERMINATION OF ANHARMONIC TERMS IN THE VIBRATIONAL HAMIL-TONIAN OF HCO^{+}.

\underline{ROBERT J. FOLTYNOWICZ}, JASON D. ROBINSON, HARTMUT G. HEDDERICH and ED-WARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

We report a harmonic frequency ω_{2} = 842.464 cm^{−1}, an anharmonic constant, x_{22} = -3.15 cm^{−1} and a vibrational angular momentum splitting constant g_{22} = 3.22 cm^{−1} for the ground state ^{1}Σ^{+} for HCO^{+}. The constants are calculated based on experimentally measured Rydberg Series from the 3pπ electronic state which converge to the rovibrational thresholds of the cation (00^{0}0), (01^{1}0), (02^{0}0) and (02^{2}0). Along with these thresholds, we use the experimentally measured x_{23} anharmonic constant from Oka et al. and the theoretical x_{12} anharmonic constant from Puzzarini et al. and Lee et al. To further characterize the bending coordinate anharmonicity of the potential energy surface of HCO^{+}, we assign Rydberg series obtained from the 3pπ ^{2}Π (030) Σ^{−} N′ = 0 and 2 converging to rovibrational thresholds (03^{1}0) and (03^{3}0) in the cation. ¿From this extrapolation, we refine our anharmonic constants for the ground electronic state for HCO^{+}.